Fermi National Laboratory


Accelerator Update for February 4, 2005 - February 7, 2005

Friday February 4
The day shift began with Operations monitoring store 3962 and the Recycler stash, with the Antiproton Source (Pbar) stacking, and MiniBooNE taking beam, and with operators just starting to send beam to the two Meson experiments MCenter and MTest.

Operations disabled all Switchyard (SY) beam to stash antiprotons in the Recycler.

Pbar resumed stacking at 3:05 PM and the Meson experiment began taking beam around 3:17 PM.

Pbar experts began a study period at 3:38 PM. They ended their study at 5:11 PM.

Saturday February 5
The midnight shift began with Operations monitoring store 3962 and the stash, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

At 12:45 AM, a Booster RF station (BBRF16) tripped off. It reset okay, but then it tripped off again at 1:11 AM and would not reset. Operators bypassed the station.

Operations terminated store 3962 at 9 AM and then put the Tevatron (TeV) into shot setup seventeen minutes later.

At 10:56 AM, the TeV suffered a sector D1 quench. Experts began their investigations. (The cause of this quench was never discovered.)

Pbar resumed stacking at 11:11 AM.

Operators resumed sending beam to MCenter and MTest at 11:13 AM.

An RF expert reported at 1:12 PM that he was cleaning the air filters for RF stations 15 & 16.

At 2:41 PM, the RF expert reported that BRF16 had a problem in the tunnel.

Operations put the TeV through a dry squeeze at 3:52 PM.

Operations reported at 6:08 PM that a Linac quadrupole power supply (QPS222) tripped off and wouldn't reset. Operators replaced all its control cards, but nothing helped. Operators called in a Linac expert.

Operations loaded a 13x0 study-store into the TeV at 9:19 PM.

Operations put the TeV into shot setup at 10 PM.

At 11:52 PM, an operator reported that he had responded to the MI62 kicker room to reset a device. Apparently, a service building door had been propped open and the operator discovered that a coyote had also come in to investigate the trip. Startled, both the operator and coyote ran. Operations eventually found the operator, but the coyote never returned, which is unfortunate since the coyote now owes Operations an investigation report. The kicker was adjusted and reset. If you see the coyote pictured below, do not approach, but please, from a safe distance, holler at him to report to the Main Control Room.

Coyote

Sunday February 6
The midnight shift began with the TeV in shot setup, and with MiniBooNE taking beam.

At 1:10 AM, a vertical separator sparked during collimator scrapping.

Operations established store 3968 at 1:17 AM with an initial luminosity of 43.11E30. The low luminosity was due to the separator spark. Pbar resumed stacking.

Operations reestablished beam to the Meson experiments at 1:32 AM.

At 3:10 AM, a Booster RF station (BRF8) tripped off and wouldn't reset. It was bypassed. Later, operators called in an expert to make repairs.

MTest experimenters reported at 5:16 AM that they were finished with beam.

Operations began stashing antiprotons into the Recycler at 5:21 AM.

A RF expert had BRF8 up and running by 9:19 AM. Pbar stacking was down for about fifteen minutes while he made his repair.

MTest asked for beam at 10:58 AM.

Operations terminated store 3968 at 8:35 PM. TeV shot setup began at 8:50 PM.

At 10:45 PM, Operations reported that the Recycler sequencer glitched and the beam went DC. There was no beam loss, but with no RF buckets to capture the beam in, there was no way to extract the antiprotons. Later, the problem was discovered. The stash will be used for the next store.

Operations established store 3969 at11:31 PM with an initial luminosity of 67.66E30. All antiprotons were loaded from the Accumulator. Pbar resumed stacking.

Operations reestablished SY beam at 11:41 PM.

Monday February 7
The midnight shift began with Operations monitoring store 3969 and the stash, with Pbar stacking, and with MiniBooNE, MTest, and MCenter taking beam.

MTest experimenters reported at 3:08 AM that they were finished with the beam.

Operations stashed antiprotons into the Recycler at 6:41 AM.

The Plan for Monday
The plan for today is to stack, stash, and store.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 2/7/2005   email Fermilab